A sample of water from a large swimming pool has a resistance of 9000 ohm at 250C when placed in a certain conductance cell. When filled with 0.02 M KCl solution, the cell has a resistance of 85 ohm at 250C. 585 gm of NaCl were dissolved in the pool, which was thoroughly stirred. A sample of this solution gave a resistance of 7500 ohm.
Given NaCl 130 ohm1cm2mol1 and KCl at 0.02M 140 ohm1cm2mol1
56. Cell constant of the conductance cell is
57. Specific conductance due to NaCl is

Grade:12

1 Answers

Arun
3 years ago
NaCl+H 
2
​ 
 O
​ 
 = 
7600
0.2346
​ 
 =3.08×10 
−5
 
NaCl
​ 
 =K 
NaCl+H 
2
​ 
 O
​ 
 −K 
2
​ 
 O
​ 
 
=(3.08−2.55)10 
−5
 
=5.3×10 
−6
 
NaCl
​ 
 = 
n
NaCl
​ 
 ×1000
​ 
 
126.5
5.3×10 
−6
 ×1000
​ 
 
=4.180×10 
−5
 
C= 
V
n
​ 
 
V= 
C
n
​ 
 = 
4.19×10 
−5
 
8.547
​ 
 =2×10 
5